OpenStudy (boxman61):

Need help squaring these.. ___ ______ √x-2 -1 = √2(x-3)

sam (.sam.):

\[\sqrt{x-2}-1=\sqrt{2(x-3)}\] \[(\sqrt{x-2}-1)^2=(\sqrt{2(x-3)})^{2}\]

sam (.sam.):

\[x-1-2\sqrt{x-2}=2(x-3)\]

sam (.sam.):

x-1-2√(x-2)=2(x-3) x-1-2√(x-2)=2x-6 -2√(x-2)=x-5 √(x-2)=(-x+5)/(2) (√(x-2))^(2)=((-x+5)/(2))^(2) x-2=((-x+5)^2)/(2)^2 x-2=((-x+5)^(2)) / 4 4x-8=(-x+5)^2 4x-8-(x^(2)-10x+25)=0 4x-8-x^(2)+10x-25=0 14x-33-x^2=0 -x^2+14x-33=0 x^(2)-14x+33=0 (x-3)(x-11)=0 x=3,11 The solution 11 is not valid, x=3

OpenStudy (boxman61):

awesome thank you.... the square was throwing me off...